sweetJ12:

points) 2. The height of one square pyramid is 24 m. A similar pyramid has a height of 8 m. The volume of the larger pyramid is 648 m3. The surface area of the smaller pyramid is 124 m^2. Determine each of the following, showing all your work and reasoning: a.) Find the scale factor that would take your from the small pyramid to the bigger one. (4 points) b.) What is the surface area of the larger pyramid? (4 points) c.) Volume of the smaller pyramid (4 points) d.) Ratio of the volume of the smaller pyramid to the larger pyramid (4 points)

jhonyy9:

any idea ? step by step ...

Narad:

(a) The scale factor is the ratio between the heights of the larger pyramid to the height of the smaller pyramid \[Ratio, r = 24/8=3\] (b) The ratio of the surface areas is \[r ^{2}=9\] Surface area of the larger pyramid is \[=9*124 =1116 m ^{2}\] (c)The ratio of the volumes is \[r ^{3}=3^{3}=27\] The volume of the smaller pyramid is \[V _{s}=648/27= 24m ^{3}\] (d) The ratio of the volume of the smaller pyramid to the larger is \[24/648= 1/27\]